Julio:

Today Nolan put a recliner on layaway by making a down payment of $90 and agreeing to pay $36 a month starting next month for 12 months. When will Nolan receive the recliner?

GoldPhenoix:

If it's asking when Nolan will receive the recliner, then I guess it's the moment he paid the down payment. But to calculate HOW much money he would need in total to buy it is: y = 90 + 36x, where y = Total cost and x = Amount of month So y = 90 + 36(12) = $522